Determine whether the matrix (

1/√3 1/√6 −1/√2

1/√3 −2/√6 0

1/√3 1/√6 1/√2

) is Orthogonal or not.


1
Expert's answer
2021-03-09T02:00:51-0500

Solution.

"A=\begin{pmatrix}
 \frac{1}{\sqrt{3}} & \frac{1}{\sqrt{6}} & -\frac{1}{\sqrt{2}} \\
 \frac{1}{\sqrt{3}} & -\frac{2}{\sqrt{6}} & 0 \\
 \frac{1}{\sqrt{3}} & \frac{1}{\sqrt{6}} & \frac{1}{\sqrt{2}} \\
\end{pmatrix}"


An orthogonal matrix is ​​a nondegenerate square matrix "A" (usually with real elements) such that

"AA^{T}=A^{T}A=I," where "A^{T}" is transposed matrix to matrix "A", and "I" is the identity matrix.

"\det A=-\frac{1}{3}-\frac{1}{6}-\frac{1}{3}-\frac{1}{6}=-1\neq0."


"A^{T}=\begin{pmatrix}
 \frac{1}{\sqrt{3}} & \frac{1}{\sqrt{3}} & \frac{1}{\sqrt{3}} \\
 \frac{1}{\sqrt{6}} & -\frac{2}{\sqrt{6}} & \frac{1}{\sqrt{6}} \\
 -\frac{1}{\sqrt{2}} & 0 & \frac{1}{\sqrt{2}} \\
\end{pmatrix}""AA^{T}=\begin{pmatrix}
 \frac{1}{\sqrt{3}} & \frac{1}{\sqrt{6}} & -\frac{1}{\sqrt{2}} \\
 \frac{1}{\sqrt{3}} & -\frac{2}{\sqrt{6}} & 0 \\
 \frac{1}{\sqrt{3}} & \frac{1}{\sqrt{6}} & \frac{1}{\sqrt{2}} \\
\end{pmatrix}
\cdot \begin{pmatrix}
 \frac{1}{\sqrt{3}} & \frac{1}{\sqrt{3}} & \frac{1}{\sqrt{3}} \\
 \frac{1}{\sqrt{6}} & -\frac{2}{\sqrt{6}} & \frac{1}{\sqrt{6}} \\
 -\frac{1}{\sqrt{2}} & 0 & \frac{1}{\sqrt{2}} \\
\end{pmatrix}=\newline
=\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 1 \\
\end{pmatrix}=I.""A^{T}A=I."

Answer. Yes, "A" is orthogonal matrix.